OpenStudy (anonymous):

Lindsey purchased a diamond necklace during the Jewelry Emporium's Semi-Annual Clearance Sale. If Lindsey paid $114.00 for the necklace and the original price is $190.00, how much of a discount did she receive? 50% 40% 60% 35%

OpenStudy (anonymous):

190*x=114 x=0.6 1-0.6=0.4 0.4*100%= 40%

OpenStudy (anonymous):

B

OpenStudy (anonymous):

114/190 to get the percentage